OpenStudy (anonymous):

Can someone help me simplify this.... Square root of 3e times square root of 12e. I don't have a square root sign here I'm sorry..

OpenStudy (anonymous):

\[\large \sqrt{3e}\sqrt{12e}\]

OpenStudy (anonymous):

\[\sqrt{3e} \times \sqrt{12e} =\sqrt{3e}\times \left( 2\sqrt{3e} \right) =2\times3\times e =6e\]

OpenStudy (anonymous):

what @CoCoTsoi said :)